Find two consecutive odd integers such that 5 times the first integer is 12 more than 3 times the second.
Answer by checkley71(8403)

X+(X+2) ARE THE 2 CONSECUTIVE ODD NUMBERS.
5X=3(X+2)+12
5X=3X+6+12
5X-3X=18
2X=18
X=18/2
X=9 ANSWER FOR THE SMALLER ODD NUMBER.
9+2=11 FOR THE LARGER NUMBER.
PROOF
5*9=3(11)+12
45=33+12
45=45